# Occam’s LGS: A simple approach for Language Gaussian Splatting

### v1

@@ Abstract
TL;DR: Gaussian Splatting is a widely adopted approach for 3D scene representation that offers efficient, high-quality 3D reconstruction and rendering. A major reason for the success of 3DGS is its simplicity of representing a scene with a set of Gaussians, which makes it easy to interpret and adapt. To enhance scene understanding beyond the visual representation, approaches have been developed that extend 3D Gaussian Splatting with semantic vision-language features, especially allowing for open-set tasks. In this setting, the language features of 3D Gaussian Splatting are often aggregated from multiple 2D views. Existing works address this aggregation problem using cumbersome techniques that lead to high computational cost and training time.
In this work, we show that the sophisticated techniques for language-grounded 3D Gaussian Splatting are simply unnecessary. Instead, we apply Occam’s razor to the task at hand and perform weighted multi-view feature aggregation using the weights derived from the standard rendering process, followed by a simple heuristic-based noisy Gaussian filtration. Doing so offers us state-of-the-art results with a speed-up of two orders of magnitude. We showcase our results in two commonly used benchmark datasets: LERF and 3D-OVS. Our simple approach allows us to perform reasoning directly in the language features, without any compression whatsoever. Such modeling in turn offers easy scene manipulation, unlike the existing methods – which we illustrate using an application of object insertion in the scene. Furthermore, we provide a thorough discussion regarding the significance of our contributions within the context of the current literature. Project Page: https://insait-institute.github.io/OccamLGS/
Numquam ponenda est pluralitas sine necessitate.
William of Ockham
Figure 1: Occam’s LGS performs training-free language feature aggregation in 3D by accurately modeling rendering of Gaussian Splatting, removing the need for expensive training in feature space. This improves runtime by two order of magnitudes, while achieving SOTA performance for open-set vision-language tasks. The straight-forward approach integrates language features seamlessly into the trained 3D Gaussian Splatting representation, allowing for downstream applications such as scene editing.
@@ 1 Introduction
3D semantic understanding is a fundamental capability in computer vision that enables machines to interpret and reason about the 3D world, with practical uses in robotics [27], autonomous navigation [42, 33], augmented reality, and interactive systems. While traditional 3D scene understanding [9] focuses on extracting geometric structure and fixed information from visual data, newer approaches integrate natural language to create richer and more meaningful scene interpretations. This emerging field, known as 3D language understanding [23, 13, 7, 17, 10, 41, 32, 4], aims to establish meaningful connections between linguistic expressions and 3D visual representations such as point clouds, meshes, and RGB-D images.
@@
Vision-language models like CLIP [2] and LSeg [5] have demonstrated the power of joint visual-linguistic understanding by training on vast image-text pairs, while foundation models like Segment Anything (SAM) [14] and DINO-V2 [20] have shown the potential of large-scale visual learning. Although these models have changed 2D image understanding completely, extending their capabilities to 3D domains remains challenging due to limited 3D-language paired data. Recent approaches [24, 21, 36, 30, 37, 1] have attempted to bridge this gap by leveraging pre-trained 2D vision-language features for 3D understanding.
However, a fundamental challenge emerges when incorporating these 2D language features into 3D representations: 2D vision-language models typically produce high-dimensional feature vectors to capture rich semantic information and 3D scenes often contain more than one million Gaussians. Generating high-dimensional features for each 3D element using a training-based approach quickly becomes computationally infeasible.
Current approaches attempt to address this challenge through scene-specific feature autoencoders [24] or quantization to create a compact discrete feature space [30] that makes training feasible. However, these scene-specific compression approaches fundamentally limit scalability and generalization. While effective for individual environments, they create barriers for multi-scene applications by requiring separate training for each new scene, increasing both computational and storage overhead. Moreover, these scene-specific solutions make it difficult to transfer knowledge between environments and limit the potential for open-world understanding.
@@
  • An training-free global optimization approach for Language 3D Gaussian Spatting, improving speed by two orders of magnitude.
  • A reasoning by rendering approach to avoiding expensive feature rendering by accurately modeling the Language 3D Gaussian Splatting forward process.
  • Compatibility to arbitrary language feature dimensionality, without feature compression.
@@
Table 1: Our approach challenges the necessity of complex modules to lift 2D language features to Gaussian Splats, resulting in a simple and fast, yet effective method.
@@ 2 Related works
Recent works [41, 3, 23, 34, 6] have explored leveraging 2D vision-language models such as CLIP [26], OpenSeg [5], and OVSeg [16], and foundation models such as SAM [14] to ground 2D language-aligned features in 3D space. Initial NeRF-based approaches [13, 15, 11] demonstrated the feasibility of distilling language features into neural radiance fields, though at the cost of slow training and rendering. More recent Gaussian-based methods have improved efficiency while following a common framework: extracting language-aligned feature maps from 2D views and grounding them to 3D space through multi-view consistency. LangSplat [24] employs SAM to obtain masks at different granularity, then extracts CLIP image features for these regions. LEGaussians [30] combines DINO [2] and CLIP features to obtain hybrid dense language feature maps. FMGS [44] similarly leverages CLIP features with DINO regularization for more robust semantic representation. GOI [25] leverages APE [29] to extract pixel-aligned features with fine boundaries.
However, incorporating these pre-trained features into 3D space presents unique challenges, particularly in managing their high dimensionality. High-dimensional features not only increase memory requirements but also significantly extend training time during optimization. Different approaches have been proposed to address this challenge: LangSplat [24] employs a scene-specific autoencoder for feature compression, LEGaussians [30] uses quantization to discretize semantic features, and Feature-3DGS [43] employs a feature compression scheme with lightweight decoder. FMGS [44] introduces multi-resolution encoding for efficient semantic embedding. GOI [25] utilizes a trainable feature clustering codebook to condense high-dimensional semantic features.
In contrast to these methods, our approach supports arbitrary-dimensional feature distillation without compromising training speed. Furthermore, our method preserves semantic meaning without requiring additional preprocessing steps or encoder-decoder structures for feature compression. This ensures generalization to the same extent as the underlying feature extractor and thus, enables a range of downstream applications requiring dynamic scenes or scene editing.
@@
The dominant paradigm in current methods [24, 30, 21, 15] involves augmenting each Gaussian with semantic features and optimizing them through iterative gradient descent. In addition, several approaches also explore the direct uplifting of 2D features to 3D. OpenGaussian [36] proposes an instance-level 3D-2D feature association method that links 3D points to 2D masks, leveraging the mIoU between rendered and ground truth masks. Semantic Gaussians [8] utilizes depth information to find corresponding 3D Gaussians along each pixel’s ray. PLGS [35] initializes semantic anchor points by backprojecting pixels into 3D space using camera parameters, while LUDVIG [18] aggregates multi-view features for each Gaussian by considering each pixel’s viewing direction.
While not focusing on semantic understanding, FlashSplat [28] demonstrates that 2D mask rendering is essentially a linear function of Gaussian labels, enabling closed-form optimization through linear programming. These methods all rely on back-projection approaches, finding Gaussians along the pixel’s ray direction and using depth information for correspondence. However, such back-projection can be inaccurate due to occlusions and incomplete depth information.
In contrast, our method accurately models the Gaussian Splatting’s forward rendering to lift 2D language-aligned features to 3D, while avoiding the back-projection limitations. Table 1 summarizes the key differences between our method and existing approaches. Unlike previous approaches, our method is training-free and does not require costly explicit semantic feature rendering. While maintaining comparable performance with existing methods, it efficiently completes the lifting of 512-dimensional features to 3D space in less than one minute.
@@ 3 Method
Figure 2: Overview of our method: Occam’s LGS consists of three main stages: (1) Forward rendering process of 3D Gaussian Splatting to obtain opacity $\alpha$, projected positions $x_{i}^{\prime}$ for each Gaussian and their corresponding pixels $p_{i}$ in 2D views (Sec. 3.1), (2) Weighted aggregation of multi-view semantic features using the alpha blending weights from the rendering process (Sec. 3.2), and (3) Filtering of noisy Gaussians that remain invisible throughout the rendering process (Sec. 3.3).
The following section provides an overview of our method, which consists of 3 main steps, Reasoning by Rendering, Weighted Feature Aggregation, and Filtering, as depicted in Figure 2. An overview of the flow of our method is further given in Algorithm 1.
3D Gaussian Splatting represents a scene as a collection of anisotropic Gaussians, where each Gaussian is defined by its position $\mathbf{x}\in\mathbb{R}^{3}$, covariance $\Sigma\in\mathbb{R}^{6}$, color represented by either a 3D color representation $\mathbf{c}\in\mathbb{R}^{3}$, or a viewpoint dependent representation as 48 spherical harmonics (SH) parameters, and opacity $o\in\mathbb{R}$
@@
The primitive representation of a Gaussian is defined as $G=\{\mathbf{x},\Sigma,\mathbf{c},o\}$, and let $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$ be the set of Gaussians in a scene. For rasterization, we first project each 3D Gaussian onto the image plane, resulting in a 2D Gaussian $G^{\prime}=\{\mathbf{x^{\prime}},\Sigma^{\prime},\mathbf{c},o\}$. Here, $\mathbf{x^{\prime}}\in\mathbb{R}^{2}$ represents the projected 2D coordinates and $\mathbf{\Sigma^{\prime}}\in\mathbb{R}^{2\times 2}$ is the projected covariance matrix.
@@
With the goal to lift multi-view 2D features into 3D space, we extend each Gaussian representation to include a high-dimensional semantic feature vector $\mathbf{f}\in\mathbb{R}^{d}$. Thus, each augmented 3D Gaussian $G_{i}$ learns its own semantic feature vector $\mathbf{f}$. Similar to color blending, the semantic feature at pixel $\mathbf{p}$ is computed through alpha-weighted blending
$\mathcal{F}(\mathbf{p})=\psi_{f}(\mathbf{p},\mathcal{G})=\sum_{i=1}^{R}\alpha_{i}\mathbf{f}_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (4)
@@
where $\psi_{f}$ is the feature rendering function and $\alpha_{i}$ represents the same Gaussian contribution as defined in the color rasterization.
@@
It is important to note that the rendering of semantics according to Equation (4) is expensive both in memory and computation when compared to the rasterization of color in Equation (2). This is due to the considerably higher dimensionality of the features $d>>3$, which prevents the GPU from fitting all required data in the shader L1 cache [24].
3.1 Reasoning by Forward Rendering
@@
Conventional methods for uplifting 2D features typically rely on back-projection, where 2D pixel-space language features are projected back to 3D space through ray-tracing. Instead, we leverage the forward rendering design of Gaussian Splatting to achieve more precise feature lifting from 2D to 3D.
Given a trained visual Gaussian model, let $S_{i}$ be the set of views where Gaussian $G_{i}$ is visible (lies within the view frustum), and $\mathbf{p}_{i}^{s}$ denote its projected pixel location in view $s$. During the forward rendering pass, we obtain the 2D features $\mathbf{f}_{i}^{s}$ at each projected pixel location $\mathbf{p}_{i}^{s}$. The 3D semantic feature $\mathbf{f}_{i}$ for Gaussian $G_{i}$ can then be computed as:
$\mathbf{f}_{i}=\mathcal{W}({\mathbf{f}_{i}^{s}}|{s\in S_{i}})$ (5)
where $\mathcal{W}$ is the weighted aggregation function as detailed in Section 3.2.
For an efficient implementation, the center locations of Gaussians are recorded during the tile-based rendering process. Specifically, for Gaussian $G_{i}$ with projected 2D center $\mathbf{x_{i}^{\prime}}$, its projected pixel location $\mathbf{p}_{i}$ is determined by
$\mathbf{p}_{i}=\lfloor\mathbf{x}_{i}^{\prime}\rfloor,$ (6)
[…]
@@
This forward feature collection approach offers several key advantages. First, the feature uplifting requires only a single forward pass through the views, making it computationally efficient. Second, view frustum culling and tile-based sorting naturally handle occluded Gaussians, ensuring we only process visible features. Finally, by explicitly recording projection locations during rendering, we establish direct and precise correspondence between 2D features and 3D Gaussians, eliminating the need for complex back-projection or iterative optimization methods.
3.2 Weighted Aggregation of Multi-View Features
Simple aggregation of multi-view features can introduce noise into semantic representations, particularly when some views have minimal Gaussian contributions. A Gaussian might be barely visible or heavily occluded in certain views, making their feature contributions less reliable. To address this, we consider the Gaussian’s contribution in each view during feature aggregation. For each Gaussian $G_{i}$, we leverage the contribution $\alpha_{i}^{s}$ at its directly projected pixel location $\mathbf{p}_{i}^{s}$ in view $s$ as the weighting factor. The weighted feature aggregation function can then be expressed as
$\mathbf{f}_{i}=\mathcal{W}({\mathbf{f}_{i}^{s},\alpha_{i}^{s}}|{s\in S_{i}})=\frac{\sum_{s\in S_{i}}\alpha_{i}^{s}\mathbf{f}_{i}^{s}}{\sum_{s\in S_{i}}\alpha_{i}^{s}}.$ (7)
This weighted aggregation ensures that views, where the Gaussian has stronger visibility and influence, contribute more significantly to its final semantic feature, resulting in more robust and consistent 3D feature representations.

### v2

# Occam’s LGS: An efficient approach for Language Gaussian Splatting

@@ Abstract
Gaussian Splatting is a widely adopted approach for 3D scene representation, offering efficient, high-quality reconstruction and rendering. A key reason for its success is the simplicity of representing scenes with sets of Gaussians, making it interpretable and adaptable. To enhance understanding beyond visual representation, recent approaches extend Gaussian Splatting with semantic vision-language features, enabling open-set tasks. Typically, these language features are aggregated from multiple 2D views, however, existing methods rely on cumbersome techniques, resulting in high computational costs and longer training times.
In this work, we show that the complicated pipelines for language 3D Gaussian Splatting are simply unnecessary. Instead, we follow a probabilistic formulation of Language Gaussian Splatting and apply Occam’s razor to the task at hand, leading to a highly efficient weighted multi-view feature aggregation technique. Doing so offers us state-of-the-art results with a speed-up of two orders of magnitude without any compression, allowing for easy scene manipulation. Project Page: https://insait-institute.github.io/OccamLGS/
Figure 1: Occam’s LGS performs training-free language feature aggregation in 3D by accurately modeling rendering of Gaussian Splatting, avoiding expensive training in feature space. This improves runtime by two orders of magnitude, while achieving SOTA performance for open-set vision-language tasks and allowing for downstream applications such as scene editing.
@@ 1 Introduction
3D semantic understanding is a fundamental capability in computer vision that enables machines to interpret and reason about the 3D world, with practical use in robotics [26], autonomous navigation [41, 32], augmented reality, and interactive systems. While traditional 3D scene understanding [9] focuses on extracting geometric structure and predefined information from visual data, newer approaches integrate natural language. This emerging field, known as 3D language understanding [22, 13, 7, 17, 10, 40, 31, 4], aims to establish rich and meaningful connections between linguistic expressions and 3D visual representations such as point clouds, meshes, and RGB-D images.
@@
Vision-language models like CLIP [2] and LSeg [5] have demonstrated the power of joint visual-linguistic understanding by training on vast image-text pairs, while foundation models like Segment Anything (SAM) [14] and DINO-V2 [19] have shown the potential of large-scale visual learning. Although these models have changed 2D image understanding completely, extending their capabilities to 3D domains remains challenging [23, 20, 35, 29, 36, 1]: directly learning 3D-language representations lacks paired data and 2D vision-language models typically produce expressive high-dimensional features, which poses an often infeasible computational cost when integrated in complex 3D scenes.
Current approaches address this challenge through scene-specific feature autoencoders [23] or quantization to create a compact feature space [29]. However, these scene-specific compression approaches fundamentally limit scalability and generalization. While effective for individual environments, they create barriers for multi-scene applications by requiring separate training for each scene, increasing both computational and storage overhead. Moreover, these scene-specific solutions make it difficult to transfer knowledge between environments and limit the potential for open-world understanding.
@@
  • A training-free global optimization approach for Language 3D Gaussian Spatting, improving speed by two orders of magnitude.
  • A reasoning by rendering approach to avoiding expensive feature rendering by following a probabilistic formulation of the Gaussian Splatting forward process.
  • Compatibility with uncompressed and arbitrary dimensional language features.
@@
Table 1: Our approach challenges the necessity of complex modules to lift 2D language features to Gaussian Splats, resulting in a simple and fast, yet probabilistically grounded method.
@@ 2 Related works
Recent works [40, 3, 22, 33, 6] have explored 2D vision-language models such as CLIP [25], OpenSeg [5], and OVSeg [16], and foundation models such as SAM [14] to ground 2D language-aligned features in 3D space. Initial approaches [13, 15, 11] demonstrated the feasibility of distilling language features into NeRF-based representations, though at the cost of slow training and rendering. More recent Gaussian-based methods have improved efficiency while following a common framework: extracting language-aligned feature maps from 2D views and grounding them to 3D space through multi-view optimization. LangSplat [23] employs SAM to obtain masks at different granularity, then extracts CLIP image features for these regions. LEGaussians [29] combines DINO [2] and CLIP features to obtain hybrid dense language feature maps and FMGS [43] uses a similar combination to improve robustness. GOI [24] leverages APE [28] to extract pixel-aligned features with fine boundaries.
However, incorporating the high-dimensional language features into 3D space presents unique challenges. It not only increases memory requirements but also significantly extends training time. Different approaches have been proposed to address this challenge: LangSplat [23] employs a scene-specific autoencoder for feature compression, LEGaussians [29] uses quantization of semantic features, and Feature-3DGS [42] employs feature compression with a lightweight decoder. FMGS [43] introduces multi-resolution encoding for efficient semantic embedding and GOI [24] utilizes a trainable feature clustering codebook.
In contrast to these methods, our approach supports arbitrary-dimensional uncompressed features without compromising training speed. Our method preserves semantic meaning without requiring additional preprocessing steps or encoder-decoder structures. This ensures open-vocabulary generalization to the same extent as the underlying feature extractor and enables a range of downstream applications using dynamic scenes or scene editing.
@@
While current methods [23, 29, 20, 15] focus on optimization-based feature lifting, several approaches also explore direct back-projection of 2D features to 3D space using pixel rays and depth information. OpenGaussian [35] proposes an instance-level 3D-2D feature association method that links 3D points to 2D masks. Semantic Gaussians [8] utilizes depth to find corresponding 3D Gaussians along each pixel’s ray. PLGS [34] initializes semantic anchor points by back-projecting pixels into 3D space using camera parameters. While not focusing on semantic understanding, FlashSplat [27] demonstrates that 2D mask rendering is essentially a linear function of Gaussian labels, enabling closed-form optimization through linear programming.
In contrast, our method accurately models Gaussian Splatting forward rendering to lift 2D language features to 3D, while avoiding back-projection limitations such as occlusion and incomplete depth information. Table 1 summarizes the key differences between our method and existing approaches. Unlike previous approaches, our method is training-free and does not require costly semantic feature rendering. While maintaining comparable performance with existing methods, it efficiently completes the lifting of 512-dimensional features to 3D in less than one minute.
@@ 3 Method
Figure 2: Overview of our method: Occam’s LGS consists of two main stages: (1) Forward rendering process of 3D Gaussian Splatting to obtain alpha blending weights $w$, projected positions $x_{i}^{\prime}$ for each Gaussian and their corresponding pixels $p_{i}$ in 2D views, followed by weighted aggregation of multi-view semantic features (Sec. 3.3); and (2) Filtering of noisy Gaussians that remain invisible throughout the rendering process (Sec. 3.4).
The following section provides an overview of our method, which establishes a theoretically grounded and efficient approach for uplifting 2D features to 3D space. It introduces the background on Gaussian splatting rendering, our formulation of maximum likelihood features uplifting, and its implementation via forward rendering and postprocessing by filtering. An overview of the flow of our method is visually presented in Figure 2 and further detailed in Algorithm 1.
3.1 Gaussian Splatting Rendering
3D Gaussian Splatting represents a scene as a collection of anisotropic Gaussians, where each Gaussian is defined by its position $\mathbf{x}\in\mathbb{R}^{3}$, covariance $\Sigma\in\mathbb{R}^{6}$, color represented by either a 3D color representation $\mathbf{c}\in\mathbb{R}^{3}$, or a viewpoint dependent representation as spherical harmonics (SH) parameters, and opacity $o\in\mathbb{R}$
@@ 3.1 Gaussian Splatting Rendering
The primitive representation of a Gaussian is defined as $G=\{\mathbf{x},\Sigma,\mathbf{c},o\}$, and let $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$ be the set of Gaussians in a scene. For rendering, we first project each 3D Gaussian onto the image plane, resulting in a 2D Gaussian $G^{\prime}=\{\mathbf{x^{\prime}},\Sigma^{\prime},\mathbf{c},o\}$, with $\mathbf{x^{\prime}}\in\mathbb{R}^{2}$ as the projected 2D coordinates and $\mathbf{\Sigma^{\prime}}\in\mathbb{R}^{2\times 2}$ as the projected covariance matrix.
@@
With the goal of lifting multi-view 2D features into 3D space, we extend each Gaussian representation to include a high-dimensional semantic feature vector $\mathbf{f}\in\mathbb{R}^{d}$. Thus, each augmented 3D Gaussian $G_{i}$ learns its own semantic feature $\mathbf{f_{i}}$. Similar to color blending, the semantic feature at pixel $\mathbf{p}$ is computed through alpha-weighted blending
$\mathcal{F}(\mathbf{p})=\psi_{f}(\mathbf{p},\mathcal{G})$ $=\sum_{i=1}^{R}\alpha_{i}\mathbf{f}_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (4)
$=\sum_{i=1}^{R}w_{i}\mathbf{f}_{i}$ (5)
@@
where $\psi_{f}$ is the feature rendering function and $w_{i}=\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ represents the alpha blending weight for Gaussian $i$.
@@
It is important to note that explicit rendering of semantics according to Equation (5) is expensive both in memory and computation when compared to the rendering of color in Equation (2). This is due to the considerably higher dimensionality of the features $d>>3$, which prevents the GPU from fitting all required data in the shader L1 cache [23].
3.2 Maximum Likelihood Feature Uplifting
@@
We deeply ground our approach to feature uplifting in the principles of 3DGS and probabilistic modeling. While conventional semantic feature learning methods [35, 8] rely on back-projection, where 2D pixel-space language features are projected back to 3D space through ray-tracing, we leverage the underlying mathematical structure of Gaussian Splatting’s forward rendering process to establish a more precise and theoretically sound feature lifting mechanism and relate it to forward modeling.
We formulate the feature uplifting problem as a Maximum-a-Posteriori (MAP) estimation problem. Given 2D feature maps $\mathbf{F}^{2D}$, the MAP estimator of 3D features $\mathbf{f}$ is:
$\arg\max_{\mathbf{f}}p(\mathbf{f}|\mathbf{F}^{2D})=\arg\max_{\mathbf{f}}p(\mathbf{F}^{2D}|\mathbf{f})\frac{p(\mathbf{f})}{p(\mathbf{F}^{2D})}$ (6)
For language gaussian splatting with uniformly distributed language features $p(\mathbf{f})$, which we assume for the efficient feature space of CLIP, the estimator simplifies to the Maximum Likelihood (ML) formulation $\arg\max_{\mathbf{f}}p(\mathbf{F}^{2D}|\mathbf{f})$.
Under a Gaussian noise model $\epsilon_{p,v}$ for language features and conditional independence of pixels, the rendered 2D feature at pixel $p$ in view $v$ is $\mathbf{f}_{p,v}^{2D}=\sum_{k=1}^{R}w_{k,p,v}\mathbf{f}_{k}+\epsilon_{p,v}$. This leads to the weighted aggregation formula for the 3D feature of each Gaussian
$\boxed{\mathbf{f}_{i}=\frac{\sum\limits_{s\in S_{i}}w_{i}^{s}\mathbf{f}_{i}^{s}}{\sum\limits_{s\in S_{i}}w_{i}^{s}}}$ (7)
where $S_{i}$ is the set of views where Gaussian $i$ is visible, $w_{i}^{s}$ represents the contribution (alpha blending weight) of Gaussian $i$ at its center projection in view $s$, and $\mathbf{f}_{i}^{s}$ is the 2D feature at Gaussian $i$’s center projected pixel in view s. This formula emerges naturally from our probabilistic framework and provides a solution that weights each view’s contribution according to the Gaussian’s influence in that view. Views, where a Gaussian is more prominent (with a higher alpha blending weight), contribute more significantly to its 3D feature, addressing the challenge of noisy or partially occluded perspectives. We provide complete step-by-step derivation in the supplementary material.
3.3 Implementation via Forward Rendering
Our weighted aggregation formula directly maps to the forward rendering process, enabling an efficient implementation that avoids complex back-projection or iterative optimization. Given a trained visual Gaussian model, for each Gaussian $G_{i}$ we perform the following steps:
  1. 1.
    Identify the set of views $S_{i}$ where the Gaussian $G_{i}$ is visible (within the view frustum).
  2. 2.
    Determine the Gaussian’s center projected pixel location $\mathbf{p}_{i}^{s}=\lfloor\mathbf{x}_{i}^{\prime}\rfloor$ in each view $s$.
  3. 3.
    Record the alpha blending weight $w_{i}^{s}$ and extract the 2D feature $\mathbf{f}_{i}^{s}$ at each center projected pixel.
  4. 4.
    Compute the weighted average following Equation 7.
@@ 3.3 Implementation via Forward Rendering
This implementation leverages the existing tile-based rendering architecture, allowing us to record projection locations and alpha values during the optimized standard rendering pass. The forward feature collection approach offers several key advantages; First, the feature uplifting requires only a single forward pass through the views, making it computationally efficient. Second, view frustum culling and tile-based sorting naturally handle occluded Gaussians, ensuring we only process visible features. Finally, by explicitly recording projection locations during rendering, we establish direct and precise correspondence between 2D features and 3D Gaussians, eliminating the need for costly back-projection or iterative optimization methods.